A sampling scheme for mechanical components from a production line calls for random samples, each consisting of eight components. Each component is classified as either good or defective. The results of 50 such samples are summa- rized in the table below. Number of Defectives Observed Frequency 30 17 From these data estimate the probability that a single component will be defective, Calculate the probabilities of various numbers of defectives in a sample of eight components, and prepare a table to compare predicted probabili- ties according to the binomial distribution with observed relative frequencies for various numbers of defectives in a sample.


1
Expert's answer
2021-11-01T13:11:18-0400

Let x= number of defectives then "X~Binomial (8,p)"

Here p is unknown, we estimate p using method of moment i.e.

"p= \frac{\sum xf}{50}=\frac{17*1+3*2}{50}=0.46"

Hence "P(X = x) = \begin{pmatrix}
 8 \\
 x
\end{pmatrix}(0.46)^x (1 - 0.46)^{8-x} ; x = 0, 1, 2, 3, ..., 8"

"P(X = 0) = (1 — 0.46)^8 = 0.0072\\
P(X = 1) = 8 * (0.46) * (1 — 0.46)^7 = 0.0493\\
P(X = 2) = (8 * 7/2) * (0.46)^2 * (1 — 0.46)^6 = 0.1469\\
P(X > 2) = 1 — P(X = 0) — P(X = 1) — P(X = 2) = 0.7966"

Expected frequencies :

Eo = 50 * P(X = 0) = 50 * 0.0072 = 0.36

E1 = 50 * P(X = 1) = 50 * 0.0493 = 2.465

E2 = 50 * P(X = 2) = 50 * 0.1469 = 7.345

E>2 = 50 * P(X > 2) = 50 * 0.7966 = 39.83